OpenStudy (anonymous):

solving linear equations : 3+t/4 = 6+1/2 ;and 2x-(5+x)/2 = 4(x+2)

OpenStudy (anonymous):

For the first linear equations multiply both the left and right hand side by 4 so that you don't have to work with fractions: 12 + t = 24 + 2 12 + t = 26...subtract 12 from both sides: t = 14

OpenStudy (anonymous):

The same concept goes for the second equation multiply both sides by the LCD: least common denominator which is 2: 4x - (5 + x) = 8(x + 2) 4x - 5 - x = 8x + 16 3x - 5 = 8x + 16 5x = 21 x = 21/5

OpenStudy (anonymous):

Actually 5x = -21 x = -21/5

OpenStudy (anonymous):

okay I am little lost on the first question, I solved down to 12+t=24 but where did 12+t=26 came from? sorry ;P

OpenStudy (anonymous):

Okay so might have forgotten the 1/2 on the right hand side when multiplied by 4 that gives you 2. That's the 2 I added to 24. so you should have: 12 + t = 24 + 2

OpenStudy (anonymous):

okay, that's great thanks!

OpenStudy (anonymous):

you're welcome